I am working two-step equations. I understand how to do all but the problems with the fractions in them. Here are two examples. Please tell me if I am doing them right
10=-30 -c/5
10=-30 -c/5
10=+c-150/5
5*10=-c-150/5 * 5/1
150+20= -c-150+150
200=-c

a/5-2=-7
a/5-2+2=-7
a/-5*5/1=-5*-5
a*25

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(75887)

10=-30 -c/5
Add 30 to both sides to get:
40=-c/5
Multiply both sides by 5 to get:
200=-c
Multiply both sides by -1 to get:
c=-200
-------------
a/5-2=-7
Add 2 to both sides to get:
a/5=-5
Multiply both sides by 5 to get;
a=-25
---------
Cheers,
Stan H.

Answer by tutorcecilia(2152)

10=-30 -c/5
10+30=-30+30 - (c/5)
40=-c/5
(5)(40) = (5)(-c/5)
200=-c
200/-1=-c/-1
-200=c
.
Check by plugging (-200=c) back into the original equation.
10=-30 -(-200)/5
10=-30 -(-200)/5
10=-30+40
10=10 [checks out]
.
.
a/5-2=-7
a/5-2+2=-7+2
a/5=-5
a/5=-5
(5)(a/5)=(5)(-5)
a=-25
.
Check by plugging (a=-25) back into the original equation.
a/5-2=-7
(-25)/5-2=-7
-5-2=-7
-7=-7 [Checks out]